Mathematics 133 – Intensive Calculus for Science 1 Quiz 6 – October 21, 2005

Your Name:

Do all work on this sheet (front and back if needed). You have 15 minutes. Calculators allowed on this quiz.

A) (6 points) Compute f′(x) for f(x) = 5x^{3} + \frac{2}{\sqrt{x}} + 2.

B) (5 points) Mark the points on the following graph where f appears not to be differentiable:

C) (9 points) The following graph shows y = f(x). At each of the marked points A, B, C, determine the signs (+, −, or 0) of f, f′, f′′ and fill in the table below: